OpenStudy (anonymous):

solve the equation -3f - 14 =1

OpenStudy (anonymous):

f=-5 I used a calculator

OpenStudy (theviper):

\[\Large{-3f-14=1}\] =>-3f=14+1=15 =>f=15/-3=-5.

OpenStudy (theviper):

clear??